40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14030     Entry time: Thu Jun 28 11:05:48 2018
Author: shruti 
Type: Update 
Category: PEM 
Subject: Seismometer temp control equipment 

Earlier today I cleared up most of the equipment at the X end near the seismometer to make the area walkable. 

In the process, I removed the connections to the temperature sensor and placed the wires on top of the can.

ELOG V3.1.3-